# Fidelity Isn’t Accuracy: When Linearly Decodable Functions Fail to Match the Ground Truth

### v1

@@ Abstract
Neural networks excel as function approximators, but their complexity often obscures the nature of the functions they learn. In this work, we propose the linearity score $\lambda(f)$, a simple and interpretable diagnostic that quantifies how well a regression network’s output can be mimicked by a linear model. Defined as the $R^{2}$ between the network’s predictions and those of a trained linear surrogate, $\lambda(f)$ offers insight into the linear decodability of the learned function. We evaluate this framework on both synthetic ($y=x\cdot\sin(x)+\epsilon$) and real-world datasets (Medical Insurance, Concrete, California Housing), using dataset-specific networks and surrogates. Our findings show that while high $\lambda(f)$ scores indicate strong linear alignment, they do not necessarily imply predictive accuracy with respect to the ground truth. This underscores both the promise and the limitations of using linear surrogates to understand nonlinear model behavior—particularly in high-stakes regression tasks.
@@ 1 Introduction
Neural networks have revolutionized supervised learning, but their internal complexity often renders their learned functions opaque. In effect, they behave as black boxes. This lack of interpretability hinders trust, accountability, and transparency—especially in high-stakes domains such as healthcare, finance, and criminal justice, where understanding a model’s decisions can be as important as their accuracy. While many techniques exist for interpreting classification networks—such as saliency maps, feature attribution methods like LIME [10] and SHAP [6], and probing approaches [1]—the interpretability of regression networks remains comparatively underexplored, especially at the level of input-output function behavior [6, 14].
One common strategy in model interpretability is to approximate a complex system with a simpler, interpretable surrogate, such as a linear model or decision tree [3, 12]. In classification settings, surrogate models are often applied to internal representations to evaluate the linear accessibility of information at each layer [1, 5]. However, in regression tasks, there has been limited work focused on assessing the linearity of a network’s full input–output function.
In this paper, we introduce a simple yet powerful idea: we quantify how linearly decodable a trained regression network’s output function is. We define a metric, $\lambda(f)$, as the coefficient of determination $R^{2}$ between a neural network’s predictions and those of a linear model trained to mimic it. We refer to this as a measure of how “linearly decodable” the network’s function is—that is, how well its input-output behavior can be approximated by a linear model. This metric offers a scalar summary of how faithfully a linear surrogate can approximate the network’s behavior.
To evaluate the utility and limitations of this approach, we perform experiments on synthetic and real-world regression datasets. For each, we compare a baseline linear model, a trained neural network, and a linear surrogate trained to mimic the network. Our findings show that $\lambda(f)$ provides a reliable measure of how linearly decodable a network’s function is—but also reveal that high fidelity to the network does not always translate into predictive success on the original task.
@@ 2 Related Works
A common approach to interpreting neural networks involves training simpler surrogate models—such as linear regressors or decision trees—to approximate their behavior. Notably, Craven and Shavlik [3] introduced TREPAN, an algorithm that uses decision trees to extract disjunctive normal form (DNF) rules by approximating the overall function computed by a neural network. Sato and Tsukimoto [12] proposed CRED, which also employs decision trees but adopts a structural approach: it trains one decision tree to approximate the output layer and additional trees for each neuron in the single hidden layer. While TREPAN operates purely on input-output behavior, CRED utilizes intermediate activations to produce layer-wise symbolic rules. In contrast to both of these algorithms, our method uses a linear surrogate not to extract rules, but to quantify the degree to which a neural network’s output function is linearly decodable.
@@
Another strategy for interpreting neural networks is probing. In this approach, smaller linear models—called “probes”—are trained to predict task-relevant labels from internal representations, providing insight into a model’s training progress or the structure of its learned features. For example, Alain and Bengio [1] introduced linear classifier probes that are inserted into each layer of a trained (or partially trained) neural network to assess how much task-relevant information is linearly accessible at that layer. This method allows researchers to analyze the flow of information during training and characterize the representational capacity of different layers. Hewitt and Manning [5] propose a structural probe to test whether syntactic information is linearly encoded in contextual word embeddings. They learn a linear transformation such that squared distances between transformed word vectors approximate syntactic distances in a dependency parse tree. This approach evaluates whether hierarchical structure is recoverable through linear geometry. While probing methods aim to characterize internal representations, our method treats the network as a black box and applies linear approximation to its overall function rather than examining representations at individual layers.
@@
Most probing methods train predictive models on frozen representations, but Saphra and Lopez [11] take a different approach. They use Singular Vector Canonical Correlation Analysis (SVCCA) [9] to track how linguistic structure emerges over the course of language model training, comparing internal representations to those learned by supervised taggers. Though not a probe in the strict sense, their work shares our goal of tracing when and how interpretable structure emerges. Our method differs by focusing on a concrete, output-level metric of linearity
@@
Other work has analyzed the complexity of neural network functions using measures such as sensitivity, Lipschitz continuity, and generalization bounds [4, 7], offering theoretical insight into expressiveness and stability. Our approach is complementary: rather than analyzing complexity directly or probing internal representations, we treat the network as a black box and ask how much of its behavior is recoverable through linear approximation. By focusing on the output function itself—rather than intermediate layers—we provide a simple, scalable diagnostic that captures one interpretable dimension of the network’s learned behavior.
@@ 3 Methodology
To assess whether a trained neural network’s learned function is linearly structured, we propose a projection-based framework that measures how closely its output can be approximated by a linear model. Our goal is not to probe internal layers or compute gradient attributions, but to treat the network as a black-box function $f\colon\mathbb{R}^{n}\to\mathbb{R}$, and ask: to what extent is the behavior of $f$ linearly decodable from its input?
We formalize this question by projecting $f$ onto the space of linear functions and measuring the quality of that approximation.
Let $\mathcal{D}\subset\mathbb{R}^{n}$ be a domain over which the network operates, and assume $f\in L^{2}(\mathcal{D})$, the space of square-integrable functions. Let $\mathcal{L}\subset L^{2}(\mathcal{D})$ denote the subspace of affine functions. We define the optimal linear surrogate as:
$g^{*}=\arg\min_{g\in\mathcal{L}}\mathbb{E}_{x\sim\mathcal{D}}\left[(f(x)-g(x))^{2}\right]$
This is the orthogonal projection of $f$ onto $\mathcal{L}$, and represents the best linear approximation to the network’s function in mean squared error.
We then define the linearity score $\lambda(f)$ as the proportion of variance in the network’s output that is captured by this surrogate:
$\lambda(f):=R^{2}(f,g^{*})=1-\frac{\mathbb{E}[(f(x)-g(x))^{2}]}{\operatorname{Var}(f(x))}$
This score takes values in $(-\infty,1]$, though it typically lies between 0 and 1. When $\lambda(f)\approx 1$, the network’s output function lies close to a linear subspace; when $\lambda(f)\approx 0$, it deviates substantially.
It is important to clarify that $\lambda(f)$ does not measure the linearity of the original data $(x,y)$, but rather the linearity of the network’s learned function $f$ with respect to its input $x$. That is, even if the underlying data is highly nonlinear, a network may learn a function that is approximately linear—making its outputs linearly decodable. Conversely, a network may learn a highly nonlinear function even on linearly structured data.
While the surrogate model $g$ provides a closed-form approximation of $f$, it is not guaranteed to preserve the predictive accuracy of the original network. As our experiments show, even high-fidelity approximations (high $\lambda(f)$) often fail to match the neural network’s performance on the original target. This distinction is central to our analysis.
Ultimately, the primary contribution of this framework is not the surrogate model itself, but the linearity score $\lambda(f)$—a simple, interpretable diagnostic of how linearly structured a network’s output function is.
@@ 4 Experiments
To evaluate the proposed framework, we apply it to synthetic and real-world datasets, including the Medical Insurance Cost dataset [2], the Concrete Compressive Strength dataset [13], and the California Housing dataset [8].
For each dataset, we evaluate three models [^1] Experimental code is available on GitHub: https://github.com/jacksoneshbaugh/lambda-linearity-score: (1) a baseline linear regression model trained directly on the input–output pairs; (2) a neural network tailored to the dataset (e.g., multi-layer perceptrons with varying layers, neurons, and ReLU activations, trained with Adam for 50–200 epochs—see Table 1); and (3) a linear surrogate model trained to approximate the output of the neural network. This setup allows us to compute and compare the linearity score $\lambda(f)$ across diverse problem domains, and to assess how closely the neural network’s behavior can be captured by a linear approximation.
@@
Table 1: Architectures of the neural networks used in each experiment. Each sequence denotes the number of units, activation functions, and any regularization layers (e.g., Dropout). These networks were selected to balance expressivity and generalization for the respective datasets.
@@ 4.1 Synthetic
We begin with a synthetic regression task defined by the function
@@
This function exhibits a continuous, nonlinear structure, providing a testbed for analyzing the expressiveness of linear approximations to neural models. We additionally introduce Gaussian noise to simulate real-world observational error. Figure 1 illustrates a sample of the dataset.
@@ 4.2 Medical Insurance Cost
The Medical Insurance Cost dataset contains diverse profiles of individuals, including features such as age, gender, BMI, number of children, and smoking status. For each profile, the dataset records an associated insurance cost. The relationships among features are moderately nonlinear, particularly for smoking status and age, making this a useful benchmark for testing linear decodability.
@@ 4.3 Concrete Compressive Strength
The Concrete Compressive Strength dataset models a nonlinear function from eight features—including cement, water, blast furnace slag, and age of the concrete sample—to compressive strength. Each input exhibits a different degree of nonlinearity with respect to the target, making this dataset well-suited for evaluating the limitations of linear surrogates.
@@ 4.4 California Housing
The California Housing dataset includes demographic and geographic information for housing blocks from the 1990 U.S. Census. Features include median income, median house age, average number of rooms and bedrooms, population, average household size, and geographic coordinates. The target variable is median house value (in hundreds of thousands of dollars). With its moderate nonlinearity and blend of structured and unstructured features, this dataset serves as a challenging and realistic case for testing how well the function can be linearly approximated.
@@ 5.1 Results
Across all experiments, we found that the linear surrogate $g$ fails to match the neural network’s predictive accuracy on the original target $y$, even when it closely approximates the network’s output $f$. In every dataset, the surrogate underperforms the neural network in predicting $y$, despite achieving high $\lambda(f)$ in some cases. This indicates that the portions of the function that escape linear approximation—though smaller in variance—are critical for prediction.
In the Medical dataset, for instance, $\lambda(f)\approx 0.93$, yet the surrogate performs worse than the baseline linear model. This suggests that the network’s predictive gains are concentrated in its nonlinear behavior, which $g$ cannot capture. In the Concrete and California datasets, we observe similar effects: the surrogate tracks the neural network well but fails to preserve its advantage. Only in the synthetic dataset is this mismatch more extreme, with low $\lambda(f)$ and complete surrogate failure. Together, these results highlight a key limitation: linear decodability does not guarantee predictive usefulness. Below, we discuss findings—summarized in Tables 2 and 3—on each dataset in detail  [^2] Due to the stochastic nature of gradient descent and nondeterministic hardware behavior, reported values (e.g., $\lambda(f)$, $R^{2}$) may vary slightly across runs. All results shown are representative of the general trends observed across repeated trials..
@@ 5.1.1 Synthetic
In the case of the synthetic dataset, our baseline linear model fails to predict the data well, yielding $R^{2}=-0.0080$. The mimic model $g$ performs nearly identically to the baseline linear regression ($R^{2}=-0.0137$), offering no meaningful improvement despite being trained to match the network’s output. Notably, both $R^{2}$ values are less than zero, meaning they both have a worse fit than a constant predictor equal to the mean of the dataset. This failure is expected: the target function is globally nonlinear, with no regions of local linearity for a simple model to exploit.
The neural network itself performs very well, achieving an $R^{2}$ score of $0.9755$. However, the surrogate’s approximation is extremely poor, yielding $\lambda(f)=-0.0105$. This indicates that the learned function $f$ is fundamentally non-linear in input space. That is, no linear transformation $g$ is able to approximate $f$ over the data distribution—even the best available linear fit performs worse than the mean of $f$. Figure 2 visualizes the outputs of all three models.
Figure 2: Predictions of the baseline linear model, neural network, and mimic model on the synthetic dataset.

### v3

# Mechanistic Interpretability

@@ Abstract
Neural networks excel as function approximators, but their complexity often obscures the types of functions they learn, making it difficult to explain their behavior. To address this, the linearity score $\lambda(f)$ is introduced, a simple and interpretable diagnostic that quantifies how well a regression network’s output can be mimicked by a linear model. Defined as the $R^{2}$ value between the network’s predictions and those of a trained linear surrogate, $\lambda(f)$ measures linear decodability: the extent to which the network’s behavior aligns with a structurally simple model. This framework is evaluated on both synthetic and real-world datasets, using dataset-specific networks and surrogates. High $\lambda(f)$ scores reliably indicate alignment with the network’s outputs; however, they do not guarantee accuracy with respect to the ground truth. These results highlight the risk of using surrogate fidelity as a proxy for model understanding, especially in high-stakes regression tasks.
@@ 1 Introduction
Neural networks have revolutionized supervised learning, but their internal complexity often renders their learned functions opaque. From an operational perspective, such models behave as black boxes. This lack of interpretability impedes trust, accountability, and transparency. While many techniques exist for interpreting classification networks—such as saliency maps, feature attribution methods like LIME ribeiroWhyShouldTrust2016 and SHAP 10.5555/3295222.3295230 , and probing approaches alainUnderstandingIntermediateLayers2018 —the interpretability of regression networks remains comparatively underexplored, specifically at the level of input-output function behavior 10.5555/3295222.3295230 .
A common strategy in model interpretability is to approximate a complex system with a simpler, interpretable surrogate, such as a linear model or decision tree 10.5555/2998828.2998832 ; satoRuleExtractionNeural2001 . In classification settings, surrogate models are often applied to internal representations to evaluate the linear accessibility of information at each layer alainUnderstandingIntermediateLayers2018 ; hewittStructuralProbeFinding2019 .
In this paper, a simple yet powerful method is proposed for probing the functional simplicity of trained regression networks: a metric, $\lambda(f)$, is defined as the coefficient of determination $R^{2}$ between a neural network’s predictions and those of a linear model trained to mimic it. Unlike interpretability techniques that produce local explanations or feature-level importance scores, this metric offers a single, global measure of the linear decodability of the network’s function. It thus complements existing interpretability approaches by providing a coarse but principled view of the network’s functional structure.
Whereas much of mechanistic interpretability focuses on understanding internal circuits, representations, or individual neurons olahZoomIntroductionCircuits2020 ; nandaProgressMeasuresGrokking2023 , this approach treats the network as a black box and investigates the type of function it implements at the input-output level. In this sense, $\lambda(f)$ can be understood as a coarse probe of the network’s functional simplicity—analogous in spirit to probing internal representations, but applied instead to the entire learned function. A general method is proposed for probing the functional simplicity of trained neural networks—a question relevant to any domain concerned with model transparency. Rather than reverse-engineering internal mechanisms, the method characterizes the functional complexity of the learned model, providing information that can support such analyses.
The utility and limitations of this approach are evaluated using experiments on synthetic and real-world regression datasets. For each, a baseline linear model, a trained neural network, and a linear surrogate trained to mimic the network are compared. These experiments show that $\lambda(f)$ provides a reliable measure of how linearly decodable a network’s function is, but also reveal that high fidelity to the network does not always translate into predictive success on the original task.
@@ 2 Related Works
A number of different methods for interpreting neural network behavior have been proposed. One common approach to interpreting neural networks involves training simpler surrogate models—such as linear regressors or decision trees—to approximate their behavior. Notably, Craven and Shavlik 10.5555/2998828.2998832 introduced TREPAN, an algorithm that uses decision trees to extract disjunctive normal form (DNF) rules by approximating the overall function computed by a neural network. Sato and Tsukimoto satoRuleExtractionNeural2001 proposed CRED, which also employs decision trees but adopts a structural approach: it trains one decision tree to approximate the output layer and additional trees for each neuron in the single hidden layer. TREPAN operates purely on input-output behavior, and CRED utilizes intermediate activations to produce layer-wise symbolic rules. In contrast to both of these algorithms, the proposed method uses a linear surrogate to quantify the linear decodability of a neural network’s output function rather than to extract rules.
@@
Probing is an alternate strategy, using smaller linear models trained to predict task-relevant labels from internal representations, providing insight into a model’s training progress or the structure of its learned features. Alain and Bengio alainUnderstandingIntermediateLayers2018 introduced linear classifier probes that are inserted into each layer of a trained (or partially trained) neural network to assess how much task-relevant information is linearly accessible at that layer. This method allows researchers to analyze the flow of information during training and characterize the representational capacity of different layers. Hewitt and Manning hewittStructuralProbeFinding2019 propose a structural probe to test whether syntactic information is linearly encoded in contextual word embeddings. They trained this probe to learn a linear transformation such that squared distances between transformed word vectors approximate syntactic distances in a dependency parse tree, thereby evaluating whether hierarchical structure is recoverable through linear geometry. Probing methods aim to characterize internal representations; however, the proposed approach treats the network as a black box and applies linear approximation to its overall function rather than examining representations at individual layers.
@@
Most probing methods train predictive models on frozen representations, but Saphra and Lopez saphraUnderstandingLearningDynamics2019 take a different approach. They use Singular Vector Canonical Correlation Analysis (SVCCA) raghuSVCCASingularVector2017 to track how linguistic structure emerges over the course of language model training, comparing internal representations to those learned by supervised taggers. Though not a probe in the strict sense, SVCCA shares the goal of tracing when and how interpretable structure emerges. The method proposed in this work differs by focusing on a concrete, output-level metric of linearity.
@@
Other work has analyzed the complexity of neural network functions using measures such as sensitivity, Lipschitz continuity, and generalization bounds haninComplexityLinearRegions2019 ; novakSensitivityGeneralizationNeural2018 , offering theoretical insight into expressiveness and stability. The proposed approach is complementary: rather than analyzing complexity directly or probing internal representations, the network is treated as a black box and the amount of its behavior recoverable through linear approximation is determined. By focusing on the output function itself rather than intermediate layers, a simple, scalable diagnostic is provided that captures one interpretable dimension of the network’s learned behavior. The method proposed in this work is not mechanistic, but it can serve as a coarse signal of functional simplicity, which may correlate with modular or interpretable internal structure—making it a useful companion to mechanistic analysis.
@@ 3 Methodology
To assess whether a trained neural network’s learned function is linearly structured, a projection-based framework is proposed that measures how closely the neural network output can be approximated by a linear model. Rather than probing internal layers or computing gradient attributions, the network is treated as a black-box function $f\colon\mathbb{R}^{n}\to\mathbb{R}$, and the degree of linear decodability of its input-output behavior is evaluated.
Formally, $f$ is projected onto the space of linear functions, and the quality of this approximation is measured.
Let $\mathcal{D}\subset\mathbb{R}^{n}$ be a domain over which the network operates, and assume $f\in L^{2}(\mathcal{D})$, the space of square-integrable functions. Let $\mathcal{L}\subset L^{2}(\mathcal{D})$ denote the subspace of affine functions. The optimal linear surrogate $g^{*}$ is then defined in Equation 1.
$g^{*}=\arg\min_{g\in\mathcal{L}}\mathbb{E}_{x\sim\mathcal{D}}\left[(f(x)-g(x))^{2}\right]$ (1)
Equation 1 defines the orthogonal projection of $f$ onto $\mathcal{L}$, representing the best linear approximation to the network’s function in mean squared error.
The linearity score $\lambda(f)$ is defined as the proportion of variance in the network’s output that is captured by $g^{*}$:
$\lambda(f):=R^{2}(f,g^{*})=1-\frac{\mathbb{E}[(f(x)-g^{*}(x))^{2}]}{\operatorname{Var}(f(x))}$ (2)
This score takes values in $(-\infty,1]$, though values typically lie between 0 and 1. When $\lambda(f)\approx 1$, the network’s output function lies close to a linear subspace; when $\lambda(f)\approx 0$, it deviates substantially.
It is important to clarify that $\lambda(f)$ does not measure the linearity of the original data $(x,y)$, but rather the linearity of the network’s learned function $f$ with respect to its input $x$. Even if the underlying data is highly nonlinear, a network may learn a function that is approximately linear—making its outputs linearly decodable. Conversely, a network may learn a highly nonlinear function even on linearly structured data.
The optimal linear surrogate $g^{*}$ provides a closed-form approximation of $f$, but it is not guaranteed to preserve the predictive accuracy of the original network. Experimental results show that even high-fidelity approximations (high $\lambda(f)$) often fail to match the neural network’s performance on the original target. The linear surrogate enables the definition of the linearity score $\lambda(f)$, a simple and interpretable diagnostic that reveals a critical distinction: fidelity to a network’s behavior does not guarantee predictive accuracy on the original task.
@@ 4 Experiments
To evaluate the proposed framework, it is applied to synthetic and real-world datasets, including the Medical Insurance Cost dataset medicalcostkaggle , the Concrete Compressive Strength dataset yehModelingStrengthHighperformance1998 , and the California Housing dataset kelleypaceSparseSpatialAutoregressions1997 .
For each dataset, three models are evaluated [^1] Experimental code is available on GitHub: https://github.com/jacksoneshbaugh/lambda-linearity-score: (1) a baseline linear regression model trained directly on the input–output pairs; (2) a neural network tailored to the dataset (e.g., multi-layer perceptrons with varying layers, neurons, and ReLU activations, trained with Adam for 50–200 epochs—see Table 1; and (3) a linear surrogate model trained to approximate the output of the neural network. This setup allows the computation and comparison of the linearity score $\lambda(f)$ across diverse problem domains, and the assessment of how closely the neural network’s behavior can be captured by a linear approximation.
@@
Table 1: Architectures of the neural networks used in each experiment. Each sequence denotes the number of units, activation functions, and any regularization layers (e.g., Dropout). These networks were designed to balance expressivity and generalization for the respective datasets.
@@ 4.1 Synthetic
The first experiment uses a synthetic regression task defined by
@@
This function exhibits a continuous, nonlinear structure, providing a testbed for analyzing the expressiveness of linear approximations to neural models. Additionally, Gaussian noise is introduced to simulate real-world observational error. Figure 1 illustrates a sample of the dataset.
@@ 4.2 Medical Insurance Cost
The Medical Insurance Cost dataset medicalcostkaggle contains diverse profiles of individuals, including features such as age, gender, BMI, number of children, and smoking status. For each profile, the dataset records an associated insurance cost. The relationships among features are moderately nonlinear, particularly for smoking status and age, making this a useful benchmark for testing linear decodability.
@@ 4.3 Concrete Compressive Strength
The Concrete Compressive Strength dataset yehModelingStrengthHighperformance1998 models a nonlinear function from eight features—including cement, water, blast furnace slag, and age of the concrete sample—to compressive strength. Each input exhibits a different degree of nonlinearity with respect to the target, making this dataset well-suited for evaluating the limitations of linear surrogates.
@@ 4.4 California Housing
The California Housing dataset kelleypaceSparseSpatialAutoregressions1997 includes demographic and geographic information for housing blocks from the 1990 U.S. Census. Features include median income, median house age, average number of rooms and bedrooms, population, average household size, and geographic coordinates. The target variable is median house value (in hundreds of thousands of dollars). With its moderate nonlinearity and blend of structured and unstructured features, this dataset serves as a challenging and realistic case for testing how well the function can be linearly approximated.
@@ 5.1 Results
Across all experiments, the optimal linear surrogate $g^{*}$ failed to match the neural network’s predictive accuracy on the original target $y$, even when it closely approximated the network’s output $f$. Furthermore, the surrogate always underperformed the neural network in predicting $y$, despite achieving high $\lambda(f)$ in some cases. This indicates that the portions of the function that escape linear approximation—though smaller in variance—are critical for prediction.
In the Medical dataset, for instance, $\lambda(f)\approx 0.93$, yet the surrogate performs worse than the baseline linear model. This suggests that the network’s predictive gains are concentrated in its nonlinear behavior, which $g^{*}$ cannot capture. In the Concrete and California datasets, similar effects are observed: the surrogate tracks the neural network well but fails to preserve its advantage. Only in the synthetic dataset is this mismatch more extreme, with low $\lambda(f)$ and complete surrogate failure. Together, these results highlight a key limitation: linear decodability does not guarantee predictive usefulness. Results are summarized in Tables 2 and 3, and findings from each experiment are discussed in detail below [^2] Due to the stochastic nature of gradient descent and nondeterministic hardware behavior, reported values (e.g., $\lambda(f)$, $R^{2}$) may vary slightly across runs. All results shown are representative of the general trends observed across repeated trials..
@@ 5.1.1 Synthetic
In the case of the synthetic dataset, the baseline linear model fails to predict the data well, yielding $R^{2}=-0.0080$. The mimic model $g^{*}$ performs nearly identically to the baseline linear regression ($R^{2}=-0.0137$), offering no meaningful improvement despite being trained to match the network’s output. Notably, both $R^{2}$ values are less than zero, meaning they both have a worse fit than a constant predictor equal to the mean of the dataset. This failure is expected: the target function is globally nonlinear, with no regions of local linearity for a simple model to exploit.
The neural network itself performs very well, achieving an $R^{2}$ score of $0.9755$. However, the surrogate’s approximation is extremely poor, yielding $\lambda(f)=-0.0105$. This indicates that the learned function $f$ is fundamentally non-linear in the input space. That is, no linear transformation is able to approximate $f$ over the data distribution—even the best available linear fit performs worse than the mean of $f$. Figure 2 visualizes the outputs of all three models.
Figure 2: Predictions from the baseline linear model, neural network, and linear surrogate on the synthetic dataset.